OpenStudy (anonymous):

help!!! when the selling price per item on 1000 items were randomly selected, they revealed the mean selling price was $10 with a standard deviation of $2. what percent of the shares sold for $12 or more?

OpenStudy (btaylor):

Assuming you are talking about a normally distributed selling price: $12 is $2 (1 standard deviation) above the mean. By the empirical rule, 68% of the data is within 1 standard deviation of the mean. So, there is 34% of the data between $10 and $12, and 50% under $10. 100-34-50=100-84=16% of the data is over $12.